# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ 7 Conclusion
In this paper, we present a novel semi-supervised deep Sobolev regressor that allows for the simultaneous estimation of the underlying regression function and its gradient. We provide a thorough convergence rate analysis for this estimator, demonstrating the provable benefits of incorporating unlabeled data into the semi-supervised learning framework. To the best of our knowledge, these results are original contributions to the literature in the field of deep learning, thereby enhancing the theoretical understanding of semi-supervised learning and gradient penalty strategy. From an application standpoint, our approach introduces powerful new tools for nonparametric variable selection and the inverse problems in partial differential equations. Moreover, our method has demonstrated exceptional performance in various numerical examples, further validating its efficacy.
@@
We would like to highlight the generality of our method and analysis, as it can be extended to various loss functions. In our upcoming research, we have extended the Sobolev penalized strategy to encompass a wide range of statistical and machine learning tasks, such as density estimation, deconvolution, classification, and quantile regression. Furthermore, we have discovered the significant role that the semi-supervised deep Sobolev regressor plays in addressing inverse problems related to partial differential equations. There still remains some challenges that need to be addressed. For example, in Theorem 5.4, the $L^{2}(\mu_{X})$-rate ${\mathcal{O}}(n^{-\frac{2s}{d+4s}})$ of the deep Sobolev regressor does not attain the minimax optimality. Additionally, the convergence rate ${\mathcal{O}}(n^{-\frac{s}{d+4s}})$ for the derivatives is also slower than the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2(s-1)}{d+2s}})$ derived in Stone (1982). Moreover, while Corollary 6.3 establishes selection consistency, it does not provide the rate of convergence. Furthermore, an interesting avenue for future research would be to investigate deep nonparametric regression with a sparse/group sparse penalty.
@@ Appendix A Supplemental Definitions and Lemmas
Let $\nu_{X}$ be a probability measure on $\Omega$ with density function $q(x)\in W^{1,\infty}(\Omega)$. Let $u\in H^{1}(\nu_{X})$ and let $v\in H^{2}(\nu_{X})$ satisfying $\nabla v\cdot\bm{n}=0$ a.e. on $\partial\Omega$, where $\bm{n}$ is the unit normal to the boundary. Then it follows that
@@
$-(\nabla u,\nabla v)_{L^{2}(\nu_{X})}$ $=-\int_{\Omega}\nabla u\cdot\nabla vqdx$
@@
where the second equality holds from integral by part, the third equality follows from the divergence theorem, and the forth one used the assumption $\nabla v\cdot\bm{n}=0$ and the equality $\nabla(\log q)=\nabla q/q$. ∎
@@
$\exp\Big{(}\frac{\lambda}{2\sigma^{2}}{\mathbb{E}}\Big{[}\max_{1\leq j\leq N}\xi_{j}^{2}\Big{]}\Big{)}\leq{\mathbb{E}}\Big{[}\max_{1\leq j\leq N}\exp\Big{(}\frac{\lambda\xi_{j}^{2}}{2\sigma^{2}}\Big{)}\Big{]}\leq N{\mathbb{E}}\Big{[}\exp\Big{(}\frac{\lambda\xi_{1}^{2}}{2\sigma^{2}}\Big{)}\Big{]}\leq\frac{N}{\sqrt{1-\lambda}},$
where the last inequality holds from Wainwright (2019, Theorem 2.6) for each $\lambda\in[0,1)$. Letting $\lambda=1/2$ yields the desired inequality. ∎
To measure the complexity of a function class, we next introduce the empirical covering number and some associated lemmas.
Definition A.3 (Empirical covering number).
is called the $L^{p}({\mathcal{D}})$ $\delta$-covering number of ${\mathcal{F}}$.
Lemma A.4 (Anthony et al. (1999, Theorem 12.2)).
@@
Notice that making a class convex-hull increases the covering number. The covering number of the convex hull was estimated by Vaart and Wellner (1996, Section 2.6.3) and Mendelson (2001, Theorem 4.1). Here we use the results derived in Vaart and Wellner (2023, Section 2.6.3), which improve the previous bounds.
Lemma A.5 (Vaart and Wellner (2023, Corollary 2.6.12)).
Let ${\mathcal{F}}$ be a set of real functions from a domain $\Omega$ to the bounded interval $[-B,B]$. Let $\delta\in(0,1)$ and ${\mathcal{D}}=\{X_{i}\}_{i=1}^{n}\subseteq\Omega$. Then for each $n\geq\operatorname{VCdim}({\mathcal{F}})$, the following inequality holds
$\log N(\delta,\operatorname{conv}({\mathcal{F}}),L^{2}({\mathcal{D}}))\leq c\operatorname{VCdim}({\mathcal{F}})B\delta^{-1},$
@@
Lemmas A.4 and A.5 demonstrate that the metric entropy of a function class is bounded by its VC-dimension. The following lemma provides a VC-dimension bound for neural network classes with a piecewise-polynomial activation function.
Lemma A.6 (Bartlett et al. (2019, Theorem 7)).
The VC-dimension of a neural network class with piecewise-polynomial activation functions is bounded as follows
$\operatorname{VCdim}({\mathcal{N}}(L,S))\leq cLS\log(S),$
where $c>0$ is an absolute constant.
By an argument similar to Duan et al. (2022, Lemma 5.7), we derive the following lemma, which shows that the first-order derivative of a ReQU neural network can be represented by a ReQU-ReLU network. With the help of this lemma and Lemma A.6, we can obtain a VC-dimension bound for the class of derivative of ReQU networks.
@@
$yz$ $=\frac{1}{4}\Big{(}\varrho_{2}(y+z)+\varrho_{2}(-y-z)-\varrho_{2}(y-z)-\varrho_{2}(z-y)\Big{)}.$ (18)
@@
By some simple calculation, we have
@@
$=\frac{1}{2}\sum_{j\in[N_{3}]}a_{pj}^{(3)}\Big{\{}\varrho_{2}\Big{(}\varrho_{1}(f_{j}^{(2)}(x))+D_{k}f_{j}^{(2)}(x)\Big{)}+\varrho_{2}\Big{(}-\varrho_{1}(f_{j}^{(2)}(x))-D_{k}f_{j}^{(2)}(x)\Big{)}$
$\quad-\varrho_{2}\Big{(}\varrho_{1}(f_{j}^{(2)}(x))-D_{k}f_{j}^{(2)}(x)\Big{)}-\varrho_{2}\Big{(}-\varrho_{1}(f_{j}^{(2)}(x))+D_{k}f_{j}^{(2)}(x)\Big{)}\Big{\}},$
@@
$S_{3,p}^{\prime,k}$ $:=\sum_{j\in[N_{3}]:a_{pj}^{(3)}\neq 0}\Big{(}S_{2,j}+S_{2,j}^{\prime,k}+12\Big{)}\leq\sum_{j\in[N_{3}]:a_{pj}^{(3)}\neq 0}\Big{(}S_{2,j}+13S_{2,j}^{\prime,k}\Big{)}$
@@
$\quad-\varrho_{2}\Big{(}\varrho_{1}(f_{j}^{(\ell-1)}(x))-D_{k}f_{j}^{(\ell-1)}(x)\Big{)}-\varrho_{2}\Big{(}-\varrho_{1}(f_{j}^{(\ell-1)}(x))+D_{k}f_{j}^{(\ell-1)}(x)\Big{)}\Big{\}}.$
@@
$\leq\sum_{j\in[N_{\ell}]:a_{pj}^{(\ell)}\neq 0}\Big{(}S_{\ell-1,j}+13(\ell-1)S_{\ell-1,j}+12S_{\ell-1,j}\Big{)}=13S_{\ell,p},$
which deduces (20) for $\ell$. Therefore, we completes the proof. ∎
@@
Notice that both ReLU and ReQU are piecewise-polynomial activation functions. By the proof of Lemma A.6 in Bartlett et al. (2019, Theorem 7), it is apparent that the VC-dimension bounds also hold for ReQU-ReLU neural networks, which are constructed in Lemma A.7. In addition, see Duan et al. (2022, Theorem 5.1) for a complete proof of the VC-dimension bound of ReQU-ReLU networks.
@@
Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. For each $\phi\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 1}$, there exists a ReQU neural network $f$ with the depth and the number of nonzero weights no more than $d\lfloor\log_{2}N\rfloor+d$ and $C^{\prime}N^{d}$, respectively, such that
$\inf_{f\in{\mathcal{F}}}\|f-\phi\|_{C^{k}(\Omega)}\leq CN^{-(s-k)}\|\phi\|_{C^{s}(K)},\quad 0\leq k\leq\min\{s,N\},$
@@
We first approximate the target function $\phi\in C^{s}(K)$ by polynomials. According to Bagby, Bos and Levenberg (2002, Theorem 2), for each $N\in{\mathbb{N}}$, these exists a polynomial $p_{N}$ of degree at most $N$ on ${\mathbb{R}}^{d}$ such that
$\sup_{x\in K}|D^{\gamma}(\phi(x)-p_{N}(x))|\leq\frac{C}{N^{s-|\gamma|}}\sum_{|\alpha|\leq s}\sup_{x\in K}|D^{\alpha}\phi(x)|,\quad 0\leq|\gamma|\leq\min\{s,N\},$ (21)
where $C$ is a positive constant depending only on $d$, $s$ and $K$. Applying Li, Tang and Yu (2019, Theorem 3.1), one obtains that there exists a ReQU neural network $f$ with the depth $d\lfloor\log_{2}N\rfloor+d$ and nonzero weights no more than $C^{\prime}N^{d}$, such that
@@
where $C^{\prime}$ is a constant independent of $N$. Combining (21) and 22 yields
$\|f-\phi\|_{C^{k}(\Omega)}\leq\sup_{x\in K}|D^{\gamma}(f(x)-p_{N}(x))|\leq CN^{-(s-k)}\|\phi\|_{C^{s}(K)},\quad 0\leq k\leq\min\{s,N\}.$
This completes the proof. ∎
Appendix B Proofs of Results in Section 3
Proof of Lemma 3.1.
By (7) and the standard variational theory, it is sufficient to focus on the variational problem

### v2

@@ VII Conclusion
In this paper, we present a novel semi-supervised deep Sobolev regressor that allows for the simultaneous estimation of the underlying regression function and its gradient. We provide a thorough convergence rate analysis for this estimator, demonstrating the provable benefits of incorporating unlabeled data into the semi-supervised learning framework. To the best of our knowledge, these results are original contributions to the literature in the field of deep learning, thereby enhancing the theoretical understanding of semi-supervised learning and gradient penalty strategy. From an application standpoint, our approach introduces powerful new tools for nonparametric variable selection. Moreover, our method has demonstrated exceptional performance in various numerical examples, further validating its efficacy.
@@
We would like to highlight the generality of our method and analysis, as it can be extended to various loss functions. In our upcoming research, we have extended the Sobolev penalized strategy to encompass a wide range of statistical and machine learning tasks, such as density estimation, deconvolution, classification, and quantile regression. Furthermore, we have discovered the significant role that the semi-supervised deep Sobolev regressor plays in addressing inverse problems related to partial differential equations. There still remains some challenges that need to be addressed. For example, in Theorem V.4, the $L^{2}(\mu_{X})$-rate ${\mathcal{O}}(n^{-\frac{2s}{d+4s}}\log^{4}n)$ of the deep Sobolev regressor does not attain the minimax optimality. Additionally, the convergence rate ${\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ for the derivatives is also slower than the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2(s-1)}{d+2s}})$ derived in [85]. Moreover, while Corollary VI.2 establishes selection consistency, it does not provide the rate of convergence. Furthermore, an interesting avenue for future research would be to investigate deep nonparametric regression with a sparse/group sparse penalty.
@@ Appendix A Supplemental Definitions and Lemmas
In this section, we present some definitions and lemmas for preparation. We first extend Green’s formula in Lebesgue measure to general measures.
@@
Let $\nu_{X}$ be a probability measure on $\Omega$ with density function $q(x)\in W^{1,\infty}(\Omega)$. Let $u\in H^{1}(\nu_{X})$ and let $v\in H^{2}(\nu_{X})$ satisfying $\nabla v\cdot{\mathbf{n}}=0$ a.e. on $\partial\Omega$, where ${\mathbf{n}}$ is the unit normal to the boundary. Then it follows that
@@
$-(\nabla u,\nabla v)_{L^{2}(\nu_{X})}$
$=-\int_{\Omega}\nabla u\cdot\nabla vqdx$
@@
where the second equality holds from integration by parts, the third equality follows from the divergence theorem [72, Theorem 1 in Section C.2], and the forth one used the assumption $\nabla v\cdot{\mathbf{n}}=0$ and the equality $\nabla(\log q)=\nabla q/q$. ∎
We next present the maximal inequality for sub-Gaussian variables.
@@
$\exp\Big{(}\frac{\lambda}{2\sigma^{2}}{\mathbb{E}}\Big{[}\max_{1\leq j\leq N}\xi_{j}^{2}\Big{]}\Big{)}$ $\leq{\mathbb{E}}\Big{[}\max_{1\leq j\leq N}\exp\Big{(}\frac{\lambda\xi_{j}^{2}}{2\sigma^{2}}\Big{)}\Big{]}$
$\leq N{\mathbb{E}}\Big{[}\exp\Big{(}\frac{\lambda\xi_{1}^{2}}{2\sigma^{2}}\Big{)}\Big{]}\leq\frac{N}{\sqrt{1-\lambda}},$
where the last inequality holds from [91, Theorem 2.6] for each $\lambda\in[0,1)$. Letting $\lambda=1/2$ yields the desired inequality. ∎
To measure the complexity of a function class, we next introduce the Vapnik-Chervonenkis (VC) dimension and some associated lemmas.
Definition 5 (VC-dimension).
A set $\{x_{i}\}_{i=1}^{m}$ is said to be shattered by ${\mathcal{F}}$ when
|$|\{(f(x_{1}),\ldots,f(x_{m})):f\in{\mathcal{F}}\}|=2^{m}.$
The VC-dimension of ${\mathcal{F}}$, denoted $\operatorname{VCdim}({\mathcal{F}})$, is the size of the largest set that can be shattered by ${\mathcal{F}}$, that is, $\operatorname{VCdim}({\mathcal{F}})=\max\{m:\Pi_{{\mathcal{F}}}(m)=2^{m}\}$. For a class ${\mathcal{F}}$ of real-valued functions, we define $\operatorname{VCdim}({\mathcal{F}})=\operatorname{VCdim}(\operatorname{sign}({\mathcal{F}}))$.
The following lemma provides a VC-dimension bound for the empirical covering number.
Lemma A.3 ([92, Theorem 12.2]).
@@
Lemma A.3 demonstrates that the metric entropy of a function class is bounded by its VC-dimension. The following lemma provides a VC-dimension bound for a deep neural network classes with a piecewise-polynomial activation function, and with a fixed architecture, i.e., the positions of the nonzero parameters are fixed.
Lemma A.4 ([93, Theorem 7]).
Let ${\mathcal{N}}$ be a deep neural network architecture with $L$ layers and $S$ non-zero parameters. The activation function is piecewise-polynomial. Then $\operatorname{VCdim}({\mathcal{N}})\leq cLS\log(S)$, where $c>0$ is an absolute constant.
With the help of Lemmas A.3 and A.4, we can bound the metric entropy of the deep neural networks by its depth and number of nonzero parameters as the following lemma. The proof of this lemma is inspired by [36, Lemma 5].
Lemma A.5.
Let ${\mathcal{N}}\subseteq{\mathcal{N}}(L,W,S)$ be a set of deep neural networks from $\Omega$ to the bounded interval $[-B,B]$. The activation function is piecewise-polynomial. Let $\delta\in(0,1)$ and ${\mathcal{D}}=\{X_{i}\}_{i=1}^{n}\subseteq\Omega$. Then
$\log N(\delta,{\mathcal{N}},L^{2}({\mathcal{D}}))\leq cLS\log(S)\log\Big{(}\frac{nB}{\delta}\Big{)},$
@@
Proof of Lemma A.5.
Before proceeding, it follows from the technique of removal of inactive nodes [36, eq. (9)] that
${\mathcal{N}}\subseteq{\mathcal{N}}(L,W,S)={\mathcal{N}}(L,W\wedge S,S).$ (16)
For each deep neural network in ${\mathcal{N}}(L,W,S)$, the number of parameters $T$ satisfies
$T$ $:=\sum_{\ell=0}^{L}(N_{\ell}+1)N_{\ell+1}\leq(L+1)2^{-L}\prod_{\ell=0}^{L}(N_{\ell}+1)$ (17)
$\leq\prod_{\ell=0}^{L}(N_{\ell}+1)\leq(W+1)^{L+1}\leq(S+1)^{L+1},$
where the last inequality is due to (16). Then there exist $\binom{T}{s}$ combinations to pick $s$ non-zero parameters from all $T$ parameters, which yields a partition
${\mathcal{N}}^{s}$ :=$:=\Big{\{}\phi\in{\mathcal{N}}:\sum_{\ell=0}^{L}(\|A_{\ell}\|_{0}+\|b_{\ell}\|_{0})=s\Big{\}}$
=$=\big{\{}{\mathcal{N}}_{1}^{s},\ldots,{\mathcal{N}}_{m}^{s}\big{\}},\quad m_{s}=\binom{T}{s},$
where the deep neural networks in the same subset have the same positions of the non-zeros parameters. Consequently,
$N(\delta,{\mathcal{N}},L^{2}({\mathcal{D}}))$
$=\sum_{s=1}^{S}N(\delta,{\mathcal{N}}^{s},L^{2}({\mathcal{D}}))=\sum_{s=1}^{S}\sum_{i=1}^{m_{s}}N(\delta,{\mathcal{N}}_{i}^{s},L^{2}({\mathcal{D}}))$
$\leq\sum_{s=1}^{S}\sum_{i=1}^{m_{s}}\Big{(}\frac{nB}{\delta}\Big{)}^{\operatorname{VCdim}({\mathcal{N}}_{i}^{s})}\leq\sum_{s=1}^{S}\binom{T}{s}\Big{(}\frac{nB}{\delta}\Big{)}^{cLs\log(s)}$
$\leq\sum_{s=1}^{S}(S+1)^{(L+1)s}\Big{(}\frac{nB}{\delta}\Big{)}^{cLs\log(s)}$
$\leq(S+1)^{(L+1)(S+1)}\Big{(}\frac{nB}{\delta}\Big{)}^{cL(S+1)\log(S)},$
where the first inequality holds from Lemma A.3, and the second inequality follows from Lemma A.4. The third inequality used the inequality $\binom{T}{s}\leq T^{s}$ and (17). Taking logarithm on both sides of the inequality yields the desired result. ∎
By an argument similar to [81, Lemma 5.7], we derive the following lemma, which shows that the first-order derivative of a ReQU neural network can be represented by a ReQU-ReLU network. With the help of this lemma and Lemma A.5, we can bound the metric entropy of the class of derivatives of ReQU networks.
@@
and
$yz=\frac{1}{4}\Big{(}\varrho_{2}(y+z)+\varrho_{2}(-y-z)\\ -\varrho_{2}(y-z)-\varrho_{2}(z-y)\Big{)}.$ (19)
@@
By some simple calculation, we have that for each $1\leq k\leq d$,
@@
$=\frac{1}{2}\sum_{j\in[N_{3}]}a_{pj}^{(3)}\Big{\{}\varrho_{2}\Big{(}\varrho_{1}(f_{j}^{(2)}(x))+D_{k}f_{j}^{(2)}(x)\Big{)}$
$\quad+\varrho_{2}\Big{(}-\varrho_{1}(f_{j}^{(2)}(x))-D_{k}f_{j}^{(2)}(x)\Big{)}$
$\quad-\varrho_{2}\Big{(}\varrho_{1}(f_{j}^{(2)}(x))-D_{k}f_{j}^{(2)}(x)\Big{)}$
$\quad-\varrho_{2}\Big{(}-\varrho_{1}(f_{j}^{(2)}(x))+D_{k}f_{j}^{(2)}(x)\Big{)}\Big{\}},$
@@
$S_{3,p}^{\prime,k}$ $:=\sum_{j\in[N_{3}]:a_{pj}^{(3)}\neq 0}\Big{(}S_{2,j}+S_{2,j}^{\prime,k}+12\Big{)}$
$\leq\sum_{j\in[N_{3}]:a_{pj}^{(3)}\neq 0}\Big{(}S_{2,j}+13S_{2,j}^{\prime,k}\Big{)}$
@@
$\quad-\varrho_{2}\Big{(}\varrho_{1}(f_{j}^{(\ell-1)}(x))-D_{k}f_{j}^{(\ell-1)}(x)\Big{)}$
$\quad-\varrho_{2}\Big{(}-\varrho_{1}(f_{j}^{(\ell-1)}(x))+D_{k}f_{j}^{(\ell-1)}(x)\Big{)}\Big{\}}.$
@@
$\leq\sum_{j\in[N_{\ell}]:a_{pj}^{(\ell)}\neq 0}\Big{(}S_{\ell-1,j}+13(\ell-1)S_{\ell-1,j}+12S_{\ell-1,j}\Big{)}$
$=13S_{\ell,p},$
which deduces (21) for $\ell$. Therefore, we complete the proof. ∎
@@
Notice that both ReLU and ReQU are piecewise-polynomial activation functions. By the proof of Lemma A.4 in [93, Theorem 7], it is apparent that the VC-dimension bounds also hold for ReQU-ReLU neural networks, which are constructed in Lemma A.6. In addition, see [81, Theorem 5.1] for a complete proof of the VC-dimension bound of ReQU-ReLU networks.
Combining Lemmas A.5 and A.6 yields the following results.
Lemma A.7.
Let ${\mathcal{N}}\subseteq{\mathcal{N}}(L,W,S)$ be a set of deep neural networks from $\Omega$ to the bounded interval $[-B,B]$. The activation function is piecewise-polynomial. Let $\delta\in(0,1)$ and ${\mathcal{D}}=\{X_{i}\}_{i=1}^{n}\subseteq\Omega$. Then
$\log N(\delta,D_{k}{\mathcal{N}},L^{2}({\mathcal{D}}))\leq cL^{2}S\log(S)\log\Big{(}\frac{nB}{\delta}\Big{)},$
where $c>0$ is an absolute constant.
@@
Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. For each $\phi\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 1}$, there exists a ReQU neural network $f$ with the depth and the number of nonzero weights no more than $d\lfloor\log_{2}N\rfloor+d$ and $C^{\prime}N^{d}$, respectively, such that $0\leq k\leq\min\{s,N\}$,
$\inf_{f\in{\mathcal{F}}}\|f-\phi\|_{C^{k}(\Omega)}\leq CN^{-(s-k)}\|\phi\|_{C^{s}(K)},$
@@
We first approximate the target function $\phi\in C^{s}(K)$ by polynomials. According to [94, Theorem 2], for each $N\in{\mathbb{N}}$, these exists a polynomial $p_{N}$ of degree at most $N$ on ${\mathbb{R}}^{d}$ such that for $0\leq|\gamma|\leq\min\{s,N\}$,
$\sup_{x\in K}|D^{\gamma}(\phi(x)-p_{N}(x))|\\ \leq\frac{C}{N^{s-|\gamma|}}\sum_{|\alpha|\leq s}\sup_{x\in K}|D^{\alpha}\phi(x)|,$ (22)
where $C$ is a positive constant depending only on $d$, $s$ and $K$. Applying [79, Theorem 3.1], one obtains that there exists a ReQU neural network $f$ with the depth $d\lfloor\log_{2}N\rfloor+d$ and nonzero weights no more than $C^{\prime}N^{d}$, such that
@@
where $C^{\prime}$ is a constant independent of $N$. Combining (22) and 23 yields
$\|f-\phi\|_{C^{k}(\Omega)}$ $\leq\sup_{x\in K}|D^{\gamma}(f(x)-p_{N}(x))|$
$\leq CN^{-(s-k)}\|\phi\|_{C^{s}(K)},$
for each $0\leq k\leq\min\{s,N\}$. This completes the proof. ∎
Appendix B Proofs of Results in Section III
Proofs of theoretical results in Section III are shown in this section.
Proof of Lemma III.1.
By (9) and the standard variational theory [72], it is sufficient to focus on the variational problem